Документ взят из кэша поисковой машины. Адрес оригинального документа : http://www.stsci.edu/smov/notes/0228smov.html
Дата изменения: Tue Mar 11 23:59:09 1997
Дата индексирования: Thu Feb 27 23:54:17 2014
Кодировка:

Поисковые слова: http www.badastronomy.com phpbb index.php
SMOV Status & Planning Meeting 02/28/97

SMOV Status & Planning Meeting Minutes

Meeting Date: 02/28/97



STATUS REPORTS

1.  Observatory Status

	MOSES status reports:

	-PCS - nominal

	-OTA - nominal

	-EPS - nominal

	-DMS - nominal
	     - a total of 5 "re-try" failures have occurred since
	       SSr installation.

	-Thermal - nominal.



2.  WFPC2 Status - nominal
	- CCD temperature now at -88C
	- The first UV monitor data, spanning a 17-hour period,
	  has been analyzed and shows no signs of UV loss.
	  The WFPC2 focus check shows the SM at +4 microns, but
	  only 2 of the three data sets have been received.
	  An STScI meeting this afternoon will determine the need
	  for a SM move and the usual form/procedure will be
	  followed.  The special SM-move ops reurest will be
	  submitted probably early this evening, for execution
	  preferrably sometime tomorrow.

3.  NICMOS Status - nominal
	
	NICMOS science continues analyzing data from
	recent internal tests.   The parallel ops test
	data analyzed to date shows no inter-camera 
	interference.  

	At 058/1506, NICMOS went into suspend state with a 
	Detector 1 VD voltage out of limits low (along with 4 others
	as reported by MOSES SI. Reason 
	unknown, but opto isolators similar to those involved
	in STIS anomalies, are strong suspects. Ball engineers
	are investingating and will present initial results/recom-
	mendations at a 11:00 am meeting this morning.  Tbere
	is not yet any reason to deviate from current SMOV plan.
 

4.  STIS Status - CCD temperature is nominal -83 degC.  MAMA low
	voltages were turned back off last evening following a
	Ball lab test result in which the MAMA high voltage came
	on while simulating the in-orbit anomaly conditions.
	Interpretation is ambiguous since it could result from the
	test set-up, but in any case, the decision was made at this
	meeting to leave the MAMA LVPSs off until the situation
	and plan are better defined.  
 
	STIS science reports that CCD read noise is at 3.9
	electrons rms, which is good.

	Note:  Yesterday's minutes omitted a report from the STIS
	IDT that the apparent noise in the 28-volt line has been 
	traced to poor grounding in the sensor circuit, and therefore
	does not represent real fluctuations.

5.  FGS status - no activity

6.  FOC/COSTAR status - nominal - no activity



PLANNING REPORTS

1.  Status of Upcoming SMSs

          	
	- SMS058 - Currently executing.
 

  	- SMS062 - Signed off and ready for uplink.
		
   	- SMS065 - Finals are out for review

   	- SMS069 - Built and initials out for review.  This SMS has
		   been rerun after a review-promoted change to
		   NICMOS Coarse Alignment Proposal (7041).


  	- SMS072 - Built and initials out for review

	- SMS076 - Built, under internal review

	- SMS079 - being worked.  This SMS currently contains the
		   1st part of the MAMA HV turn-on, per the nominal
		   plan.


2.  Replans - none

OPERATIONS STATUS - 

1.  OPUS
	- Yesterday's STIS science-format data failed during
	  CALSTIS.  The problem has been traced to a TRANS problem
	  that miscongfigures a PMDB flag.  The problem was corrected
	  manually in the PMDB and a permanent fix will is being
	  investigated by John Baum.  The data then calibrated suc-
	  cessfully but then experienced another hang-up traced to a
	  header problem, which is under study.

	  OPUS Generic Conversion had trouble handling a 
	  Reed-Solomon missing packet error;  Jim rose is looking
	  into it this morning.

2.  Archive - Nominal



biagetti@stsci.edu

Go Back Go back to the Meeting Minutes Page